A 50 inch diameter steel pipe with a thickness of 0.25 inches, carries oil of specific gravity of 0.822 under a head of 350 ft of oil. Compute the pressure inside the pipe in psi, circumferential stress, Compute the thickness of steel required to carry a pressure of 220 psi with a hoop stress of 18,000 psi.
Answer #1

daimeter of pipe ÷12= 50÷12 =4.16 in > 0.25 in.so,pipe is thin cylinder.

specific gravity of oil = 0.822

density of oil = 0.822×1 = 0.822 g/cc

=0.822×0.036127 =0.02969 lb/in^3

pressure inside pipe p = \rho gh =0.02969×3.8622×350×12=481.608 psi

Hoop stress = pd÷2t =(481.608×50)÷(2×0.25)

= 48160.8 psi

given that pressure p =220 psi

hoop stress = 18000 psi

we know that hoop stress = pd÷2t

= (220×50)÷(2×t)=18000

t = (220×50)÷(2×18000)

thickness of pipe t = 0.3055 inches.